In a town of 500 people, 298 have bicycle, 225 have bike and 140 have car. 29 have both bicycle and bike, 35 have both bike and car, 20 have both bicycle and car and 11 have all the three. How many have only car?
Answer by ikleyn(52781)

.

    Car_only = Car - (Car and Bicycle) - (Car and Bike) + (All_the_three) = 140 - 20 - 35 + 11 = 96.    ANSWER

----------------

It is nice problem.

If you want to learn  MORE  about this subject,  look into my lessons
    - Counting elements in sub-sets of a given finite set
    - Advanced problems on counting elements in sub-sets of a given finite set
    - Challenging problems on counting elements in subsets of a given finite set
in this site.

They are really  SWEET  and  CRISPY  problems,  and  SWEET  and  CRISPY  lessons.

It is worth to read (!)

-------------

When you complete reading,  come again and ask me,  what to read next . . .